If a,b,c are in A.P. Prove that b+c,c+a,a+b are also in A.P.?

Grade:12th pass

11 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago
Dear Rohan
If a,b,c are in AP then 2b=a+c
We have to prove
2(a+c)=(b+c)+(a+b)
so 2a+2c=2b+a+c
so a+c=2b
Which is true
Hence proved

All the best.
AKASH GOYAL
AskiitiansExpert-IITD

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

In IIT JEE exams emphasis is more on the understanding of concepts rather than rote learning and memorizing of formulae. So read good book and be attentive. Do one or two book only for each subject but revise them regularly. Revision is very important to secure good marks. so don't run after too many books and material. You can use following books for the preparation.
Mathematics:- RD Sharma and Arihant Books, NCERT
Physics H. C. Verma, NCERT. D C PANDEY
Chemistry
NCERT AND
P. Bahadur - Physical Chemistry
Arihant Publications - Organic Chemistry
O.P. Agarwal - IIT Chemistry - for Inorganic Chemistry
vikas askiitian expert
13 years ago

 if a,b,c are in AP then

  b-a=c-b

   a+c=2b ...........1

now we have to prove that b+c,c+a,a+b are in AP,

     let these are in AP then

                 (a+c)-(b+c)=(a+b)-(a+c)

                      (a+c)=2b

from eq 1 we have a+c=2b ,RHS=LHS

so these terms are also in AP...

 

Poornima
7 years ago
a,b,c. Are in apSo their common difference will be equalSo,b-a=c-ba+c=2bTo proveb+c,c+a,a+b are in apProofLet us consider that they are in apSo,(C+a)-(b+c)=(a+b)-(c+a)c+a-b-c=a+b-c-aa-b=b-ca+c=2bTheir common difference is equal.So they are in ap
Ashutosh Kumar
7 years ago
Solution-a,b&c are in ap .Now , a+(-a-b-c),b+(-a-b-c)& c+(-a-b-c) are in ap.Now,-b-c,-a-c&-a-b are in apOr,-(b+c),-(a+c)&-(a+b)are in apor,b+c,a+c&a+b are in ap
Ashutosh Bajpai
7 years ago
If b+c ,c+a,a+b are in AP then c+a -(b+c)=a+b-(c+a)a-b =b-cThen 2b=a+cThus it proved that a,b and c are in Ap
vallish
7 years ago
here a b c are in ap then
b-a=c-b take – common
-(a-b)=-(b-c) divide both by -
a-b=b-c.........(1)
assume a+b b+c a+c are ap
b+c -(a+b)=a+c-(b+c)
a-b=b-c.........(2)
from 1 & 2 hence proof
Ritika Das
7 years ago
Well, since it is given that a,b and c are in Arithmetic Progression, for any arbitrary common difference d:-
b=a+d ;
c=a+2d.
Therefore, 
a+b=2a+d;
a+c=2a+2d;
b+c=2a+3d.
Hete all the expressions have a common difference d.
Thus proved that a+b, a+c and b+c are in Arithmetic Progression for the common difference d.
Prabhleen kaur
7 years ago
If a,b,c are in AP. So,b-a=c-b0=c+a-b-b0=c+a-2b2b=c+a ..................1Let us assume that b+c, c+a, a+b are in AP.So, (c+a)-(b+c) = (a+b)-(c+a)(a+c)=2b2b=2b [from eq 1]LHS=RHSTherefore, b+c, c+a and a+b are in AP. Hence,proved
Prabhleen kaur
7 years ago
A+c =2b.......1Let us assume that b+c,c+a and a+b are in AP(c+a)-(b+c)=(a+b)-(c+a)a+c=2b2b=2b[from eq 1]LHS=RHSTHEREFORE B+C,C+A AND A+B ARE IN AP.hence, proved
ARKADEEP DAS
7 years ago
let common differene be d
then b=a+d
c=a+2d
now a+c = 2a+2d
b+c=2a+3d
a+b=2a+d
so see there are in ap
ok rohan u are done
Kushagra Madhukar
4 years ago
Dear student,
Please find the solution to your problem below.
 
If a,b,c are in AP then
b – a = c – b
a + c = 2b
Adding, a + c on both sides
Hence, 2(a + c) = 2b + a + c
or, 2(a + c) = b + a + b + c
or, 2(a + c) = (a + b) + (b + c)
Hence, (b + c), (c + a), (a + b) are also in AP
 
Hope it helps.
Thanks and regards,
Kushagra